# Interactive and Hybrid Imitation Learning: Provably Beating Behavior Cloning

### v2

@@ 1 Introduction
Imitation learning, or learning from demonstrations, is a widely applied paradigm for learning sequential decision-making policies pomerleau1988alvinn; bojarski2016end; bansal2018chauffeurnet. In many applications, it offers a preferable alternative to reinforcement learning, as it bypasses the need for carefully designed reward functions and avoids costly exploration osa2018algorithmic; sun2017deeply.
Two prominent data collection regimes exist in imitation learning: offline and interactive. In offline imitation learning, expert demonstration data in the format of trajectories is collected ahead of time, which is a non-adaptive process that is easy to maintain. In contrast, in interactive imitation learning, the learner is allowed to query the expert for annotations in an adaptive manner ross2011reduction; ross2014reinforcement; sun2017deeply. The most basic and well-known approach for offline imitation learning is Behavior Cloning ross2010efficient; foster2024behavior, which casts the policy learning problem as a supervised learning problem that learns to predict expert actions from states. Although simple and easy to implement, offline imitation learning has the drawback that the quality of the data can be limited pomerleau1988alvinn. As a result, the trained model can well suffer from compounding error, where imperfect imitation leads the learned policy to enter unseen states, resulting in a compounding sequence of mistakes. In contrast, in interactive imitation learning, the learner maintains a learned policy over time, with the demonstrating experts providing corrective feedback on-policy, which enables targeted collection of demonstrations and improves sample efficiency.
Recent work foster2024behavior, via a sharp theoretical analysis of Behavior Cloning, shows that the sample efficiency of Behavior Cloning cannot be improved in general when measuring using the number of trajectories annotated. Interactive methods like DAgger ross2010efficient can enjoy sample complexity benefits, but so far the benefits are only exhibited in limited examples, with the most general ones in the tabular setting rajaraman2021value. This leaves open the question:
Can interaction provide sample efficiency benefit for imitation learning under a broad range of settings, especially with function approximation?
@@
    Motivated by the costly nature of interactive labeling on entire trajectories laskey2017comparing; mandlekar2020human, we propose to measure the cost of annotation using the number of states annotated by the demonstrating expert. We propose a general state-wise interactive imitation learning algorithm, Stagger, and show that as long as the expert can recover from mistakes at low cost in the environment ross2011reduction, it significantly improves over Behavior Cloning in terms of its number of state-wise demonstrations required.
@@
    Motivated by practical imitation learning applications where sets of offline demonstration data are readily available, we study hybrid imitation learning, where the learning agent has the additional ability to query the demonstration expert interactively to improve its performance. We design a hybrid imitation learning algorithm, Warm-Stagger, and prove that its policy optimality guarantee is not much worse than using either of the data sources alone.
@@
    Inspired by compounding error pomerleau1988alvinn and cold start problem lin2020autoregressive; oetomo2023cutting, two practical challenges in imitation learning, we provide an MDP example, for which we show hybrid imitation learning can achieve strict sample complexity savings over using either source alone, and provide simulation results that verify this theoretical claim.
@@ 2 Preliminaries
Define $n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:n}$ to represent the sequence $\{x_{i}\}_{i=1}^{n}$. We will frequently use the Hellinger distance to measure the difference between two distributions: $D_{H}^{2}(\mathbb{P},\mathbb{Q})=\int(\sqrt{\tfrac{d\mathbb{P}}{d\omega}}-\sqrt{\tfrac{d\mathbb{Q}}{d\omega}})^{2}d\omega$, where $\mathbb{P}$ and $\mathbb{Q}$ share a dominating measure $\omega$.
@@
An episodic MDP $\mathcal{M}$ is defined as a tuple $(\mathcal{S},\mathcal{A},P,\mathcal{R},H)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:=\{P_{h}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})\}_{h=1}^{H}$ denotes the transition dynamics, $\mathcal{R}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta([0,1])$ denotes the reward distribution, and $H$ denotes episode length. Given a stationary policy $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$, we use $\pi(\cdot|s)$ to denote the action distribution of $\pi$ on $s$. A policy induces a distribution over trajectories $\tau=(s_{h},a_{h},r_{h})_{h=1}^{H}$ by first drawing the initial state $s_{1}\sim P_{0}(\varnothing)$, and then iteratively taking actions $a_{h}\sim\pi(\cdot\mid s_{h})$, receiving rewards $r_{h}\sim\mathcal{R}(s_{h},a_{h})$, and transitioning to the next state $s_{h+1}\sim P_{h}(s_{h},a_{h})$ (except at step $H$, where $P_{H}=\varnothing$). Let $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ denote expectation and probability law for $(s_{h},a_{h},r_{h})_{h=1}^{H}$ induced by $\pi$ and $\mathcal{M}$. Given $\pi$, denote by $d^{\pi}(s):=\frac{1}{H}\mathbb{P}^{\pi}(s_{h}=s)$ its state visitation distribution. The expected return of policy $\pi$ is defined as $J(\pi):=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}r_{h}\right],$ and the value functions of $\pi$ are given by $V_{h}^{\pi}(s):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s\right]$, and $Q_{h}^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s,a_{h}=a\right].$ If for policy $\pi$, step $h$, and state $s$, $\pi(\cdot\mid s)$ is the delta-mass on an action, we also sometimes slightly abuse the notation and let $\pi(s)$ denote that action.
@@
Throughout, we assume the access to an Markovian policy class $\mathcal{B}$ of finite size $B$, which contains the deterministic expert demonstrator policy $\pi^{\mathrm{E}}:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$. A $(\text{MDP},\text{Expert})$ pair $(\mathcal{M},\pi^{\mathrm{E}})$ is said to be $\mu$-recoverable if for all $h\in[H]$, $s\in{\mathcal{S}}$ and $a\in\mathcal{A}$, $Q_{h}^{\pi^{\mathrm{E}}}(s,a)-V_{h}^{\pi^{\mathrm{E}}}(s)\leq\mu$. Additionally, we assume normalized return foster2024behavior, where for any trajectory $(s_{h},a_{h},r_{h})_{h=1}^{H}$, $\sum_{h=1}^{H}r_{h}\in[0,R]$. Throughout this paper, we make the assumption that our imitation learning problem is deterministically realizable:
@@
Definition 1 (Each-step Mixing of $\mathcal{B}$).
$\bar{\Pi}_{\mathcal{B}}=\{\bar{\pi}_{u,h}(a|s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a|s):u\in\Delta(\mathcal{B})\}.$
An each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and takes action $a\sim\pi(\cdot|s)$ (e.g. li2022efficient; li2023agnostic). Observe that $\bar{\pi}_{u}$ is a Markovian policy.
@@
In offline imitation learning, the agent is given a collection of expert trajectories $\mathcal{D}=\left\{\tau_{1},\ldots,\tau_{N_{\mathrm{off}}}\right\}$, where $\tau_{i}=(s_{i,h},a_{i,h})_{h=1}^{H}$ is the $i$-th (reward-free) trajectory, all of which are drawn iid from the trajectory distribution of expert policy $\pi^{\mathrm{E}}$. Behavior Cloning finds a policy $\pi\in\mathcal{B}$ that minimizes its log loss on expert’s actions on the seen states, i.e.,
@@
Recent result of foster2024behavior establishes a horizon-independent analysis of Behavior Cloning, which we recall its guarantees here:
Theorem 2 (Guarantee of BC foster2024behavior).
@@
In interactive IL, the learner has the ability to query the demonstration expert interactively. A first way to model interaction with expert is through a trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ ross2011reduction; foster2024behavior: given a state sequence $(s_{h})_{h=1}^{H}$, return $(a_{h})_{h=1}^{H}$ such that $a_{h}=\pi^{\mathrm{E}}(s_{h})$ for all $h$. Subsequent works have considered modeling the expert as a state-wise demonstration oracle judah2014active; brantley2019disagreement; nguyen2020active; sekhari2024selective $\mathcal{O}^{\mathrm{State}}$: given a state $s_{h}$ and step $h$, return $a_{h}=\pi^{\mathrm{E}}(s_{h})$. We consider the learner interacting with the environment and demonstration oracles using the following protocol:
@@
In practice, we expect the cost of querying $\mathcal{O}^{\mathrm{Traj}}$ to be higher than that of collecting a single offline expert trajectory laskey2017comparing. Since $H$ queries to $\mathcal{O}^{\mathrm{State}}$ can simulate one query to $\mathcal{O}^{\mathrm{Traj}}$, the cost of a single $\mathcal{O}^{\mathrm{State}}$ query should be at least $\frac{1}{H}$ the cost of $\mathcal{O}^{\mathrm{Traj}}$. Consequently, we also expect one $\mathcal{O}^{\mathrm{State}}$ query to be more expensive than obtaining an additional offline (state, expert action) pair. We denote the ratio between these two costs as $C$, where $C\geq 1$ is an application-dependent constant.  [^1] For practical settings such as human-in-the-loop learning with expert interventions mandlekar2020human; spencer2020learning, obtaining a short segment of corrective demonstrations may be cheaper than querying $\mathcal{O}^{\mathrm{State}}$ for each state therein. Here, we focus on a simplified setting and leave detailed cost modeling for such settings as interesting future work.
@@ 3 State-wise Annotation in Interactive Imitation Learning
Recent work foster2024behavior on refined analysis of Behavior Cloning (BC) casts doubt on the utility of interaction in imitation learning: when measuring sample complexity in the number of trajectories annotated, BC is minimax optimal even among interactive algorithms (foster2024behavior, Corollary 2.1 and Theorem 2.2). Although benefits of interactive approaches have been shown in specific examples, progresses so far have been sparse foster2024behavior; rajaraman2021value, with the most general results in the less-practical tabular setting rajaraman2021value. In this section, we show that interaction benefits imitation learning in a fairly general sense: when measuring sample complexity using the number of state-wise annotations, we design an interactive algorithm with sample complexity better than BC, as long as the expert policy has a low recovering cost $\mu$ in the environment.
@@
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$ with query budget $N_{\mathrm{int}}$, Markovian policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$.
@@ 3.1 Interactive IL Enables Improved Sample Complexity with State-wise Annotations
Our algorithm Stagger (short for State-wise DAgger), namely Algorithm 1, interacts with the demonstration expert using a state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$. Similar to the original DAgger ross2011reduction, it requires base policy class $\mathcal{B}$ and reduces interactive imitation learning to no-regret online learning. At round $n$, it rolls out the current policy $\pi^{n}$ obtained from an online learning oracle $\mathbb{A}$ and samples $s^{n}$ from $d^{\pi^{n}}$. A classical example of $\mathbb{A}$ is the exponential weight algorithm that chooses policies from $\bar{\Pi}_{\mathcal{B}}$ (cesa2006prediction; see Proposition 38 in Appendix F). It then queries $\mathcal{O}^{\mathrm{State}}$ to get expert action $a^{*,n}$ and updates $\mathbb{A}$ with loss function $\ell^{n}(\pi)$ induced by this new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}^{N_{\mathrm{int}}}_{n=1}$, i.e., sample one $\pi^{n}$ uniformly at random and execute it for the episode. In contrast to the DAgger variant analyzed in foster2024behavior, which trains a distinct policy at each step—yielding $H$ policies in total—and employs trajectory-level annotations, our algorithm utilizes parameter sharing and uses state-wise annotations.
@@
Suppose Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with decision space $\Delta(\mathcal{B})$ and returns each-step mixing policies $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$. Then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
@@
Theorem 3 shows that Stagger returns a policy of suboptimality $O(\frac{\mu H\log B}{N_{\mathrm{int}}})$ using $N_{\mathrm{int}}$ interactive state-wise annotations from the expert. In comparison, with the cost of $N_{\mathrm{int}}$ state-wise annotations, one can obtain $\frac{CN_{\mathrm{int}}}{H}$ trajectory-wise annotations; foster2024behavior’s analysis shows that Behavior Cloning with this number of trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\log B}{CN_{\mathrm{int}}})$; recall Theorem 2. Thus, if $C\ll\frac{R}{\mu}$, Algorithm 1 has a better cost-efficiency guarantee than Behavior Cloning.
We now sketch the proof of Theorem 3. In line with foster2024behavior, we define the online, on-policy state-wise estimation error as
@@
The proof proceeds by bounding this error and translating it to the performance difference between $\hat{\pi}$ and $\pi^{\mathrm{E}}$. While our definition of estimation error is similar to (foster2024behavior, Appendix C.2), their definition requires all $H$ states per trajectory, while ours depends on the distribution over a state sampled uniformly from the rollout of policy $\pi^{n}$. This enables each labeled state to serve as immediate online feedback, fully utilizing the adaptivity of online learning. In contrast, trajectory-wise annotations may cause the online learning oracle to operate under delayed feedback joulani2013online; wan2022online, which incurs a fundamental extra factor of $H$ in state-wise sample complexity compared to our approach.

### v3

@@ 1 Introduction
Imitation learning, also known as learning from demonstrations, is a widely-used approach for learning policies to make sequential decisions Pomerleau (1988); Bojarski et al. (2016); Bansal et al. (2018). In many applications, it offers a preferable alternative to reinforcement learning, as it bypasses the need for carefully designed reward functions and avoids costly exploration Osa et al. (2018); Sun et al. (2017).
Two prominent data collection regimes exist in imitation learning: offline and interactive. In offline imitation learning, expert demonstration data in the format of trajectories is collected ahead of time, which is a non-adaptive process that is easy to maintain. In contrast, in interactive imitation learning, the learner is allowed to query the expert for annotations in an adaptive manner Ross et al. (2011); Ross and Bagnell (2014); Sun et al. (2017). The most basic and well-known approach for offline imitation learning is Behavior Cloning Ross and Bagnell (2010); Foster et al. (2024), which casts the policy learning problem as a supervised learning problem that learns to predict expert actions from states. Although simple and easy to implement, offline imitation learning has the drawback that the quality of the data can be limited Pomerleau (1988). As a result, the trained model can well suffer from compounding error, where imperfect imitation leads the learned policy to enter unseen states, resulting in a compounding sequence of mistakes. In contrast, in interactive imitation learning, the learner maintains a learned policy over time, with the demonstrating experts providing corrective feedback on-policy, which enables targeted collection of demonstrations and improves sample efficiency.
Recent work Foster et al. (2024), via a sharp analysis of Behavior Cloning, shows that the sample efficiency of Behavior Cloning cannot be improved in general when measuring using the number of trajectories annotated. Interactive methods like DAgger Ross and Bagnell (2010) can enjoy sample complexity benefits, but so far the benefits are only exhibited in limited examples, with the most general ones in the tabular setting Rajaraman et al. (2021). This leaves open the question:
Can interaction provide sample efficiency benefit for imitation learning under diverse settings, notably in the presence of function approximation?
@@
    Motivated by the costly nature of interactive labeling on entire trajectories Laskey et al. (2017a); Mandlekar et al. (2020), we propose to measure the cost of annotation using the number of states annotated by the demonstrating expert. We propose a general state-wise interactive imitation learning algorithm, Stagger, and show that as long as the expert can recover from mistakes at low cost in the environment Ross et al. (2011), it significantly improves over Behavior Cloning in terms of its number of state-wise demonstrations required.
@@
    Motivated by practical imitation learning applications where sets of offline demonstration data are readily available, we study hybrid imitation learning, where the learning agent can additionally query the demonstration expert interactively to improve its performance. We design a hybrid imitation learning algorithm, Warm-Stagger, and prove that its policy optimality guarantee is not much worse than using either of the data sources alone.
@@
    Inspired by compounding error Pomerleau (1988) and cold start problem Lin et al. (2020); Oetomo et al. (2023), we provide an MDP example, for which we show hybrid imitation learning can achieve strict sample complexity savings over using either source alone, and provide simulations that verify this claim.
@@ 2 Preliminaries
Define $n]:=\left\{1,\ldots,n\right\}$. Denote by $\Delta(\mathcal{X})$ the set of probability distributions over a set $\mathcal{X}$. For $u\in\Delta(\mathcal{X})$ and $x\in\mathcal{X}$, we denote by $u(x)$ the $x$-th coordinate of $u$ and $e_{x}$ the delta mass on $x$. We use the shorthand $x_{1:n}$ to represent the sequence $(x_{i})_{i=1}^{n}$; we will also apply this shorthand to tuples, e.g. using $(x,y)_{1:n}$ to denote $(x_{i},y_{i})_{i=1}^{n}$. We will frequently use the Hellinger distance to measure the difference between two distributions: $D_{H}^{2}(\mathbb{P},\mathbb{Q})=\int(\sqrt{\tfrac{d\mathbb{P}}{d\omega}}-\sqrt{\tfrac{d\mathbb{Q}}{d\omega}})^{2}d\omega$, where $\mathbb{P}$ and $\mathbb{Q}$ share a dominating measure $\omega$.
@@
A fixed-horizon episodic MDP $\mathcal{M}$ is defined as a tuple $(\mathcal{S},\mathcal{A},P,\rho,\mathcal{R},H)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})$ denotes the transition dynamics, $\rho\in\Delta({\mathcal{S}})$ is the initial state distribution, $\mathcal{R}:{\mathcal{S}}\times\mathcal{A}\rightarrow\Delta([0,1])$ denotes the reward distribution, and $H$ denotes episode length.  [^1] Here we assume that the transition dynamics and reward functions are stationary, i.e., it does not depend on time step in the episode. To translate our results to the nonstationary transition dynamics and reward setting (to have a fair comparison with (e.g., Foster et al., 2024; Rajaraman et al., 2020)), we can augment the state with a step index, i.e. define $\tilde{s}=(s,h)$. Given a stationary policy $\pi:{\mathcal{S}}\rightarrow\Delta(\mathcal{A})$, we use $\pi(\cdot|s)$ to denote the action distribution of $\pi$ on state $s$. Rolling out policy $\pi$ in MDP $\mathcal{M}$ gives a distribution over trajectories $\tau=(s_{h},a_{h},r_{h})_{h=1}^{H}$ by first drawing the initial state $s_{1}\sim\rho$, and then iteratively taking actions $a_{h}\sim\pi(\cdot\mid s_{h})$, receiving rewards $r_{h}\sim\mathcal{R}(s_{h},a_{h})$, and transitioning to the next state $s_{h+1}\sim P(s_{h},a_{h})$ (except at step $H$). Let $\mathbb{E}^{\pi}$ and $\mathbb{P}^{\pi}$ denote expectation and probability law for $(s_{h},a_{h},r_{h})_{h=1}^{H}$ induced by $\pi$ and $\mathcal{M}$. Given $\pi$, denote by $d^{\pi}(s):=\frac{1}{H}\sum_{h=1}^{H}\mathbb{P}^{\pi}(s_{h}=s)$ its state visitation distribution. The expected return of policy $\pi$ is defined as $J(\pi):=\mathbb{E}^{\pi}\left[\sum_{h=1}^{H}r_{h}\right],$ and the value functions of $\pi$ are given by $V_{h}^{\pi}(s):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s\right]$, and $Q_{h}^{\pi}(s,a):=\mathbb{E}^{\pi}\left[\sum_{h^{\prime}=h}^{H}r_{h^{\prime}}\,|\,s_{h}=s,a_{h}=a\right].$ If for policy $\pi$, step $h$, and state $s$, $\pi(\cdot\mid s)$ is the delta-mass on an action, we also sometimes slightly abuse the notation and let $\pi(s)$ denote that action.
@@
Throughout, we assume the access to a class $\mathcal{B}$ of stationary policies of finite size $B$. A $(\text{MDP},\text{Expert})$ pair $(\mathcal{M},\pi^{\mathrm{E}})$ is said to be $\mu$-recoverable if for all $h\in[H]$, $s\in{\mathcal{S}}$ and $a\in\mathcal{A}$, $Q_{h}^{\pi^{\mathrm{E}}}(s,a)-V_{h}^{\pi^{\mathrm{E}}}(s)\leq\mu$. Additionally, we assume normalized return Foster et al. (2024), where there exists some $R>0$ such that for any trajectory $(s_{h},a_{h},r_{h})_{h=1}^{H}$ rolled out in $\mathcal{M}$, $\sum_{h=1}^{H}r_{h}\in[0,R]$. It always holds that $\mu\leq R$, while in many applications we expect $\mu$ to be much smaller (Ross et al., 2011, Section 2.2). Throughout this paper, we make the assumption that our imitation learning problem is deterministically realizable:
@@
Definition 1 (Each-step mixing and each-step mixing policy class).
For $u$ in $\Delta(\mathcal{B})$, define its each-step mixing policy $\bar{\pi}_{u}$ as a stationary policy from $\mathcal{S}$} to $\Delta(\mathcal{A})$:
$\bar{\pi}_{u}(a\mid s):=\sum_{\pi\in\mathcal{B}}u(\pi)\pi(a\mid s).$
Define $\bar{\Pi}_{\mathcal{B}}:=\left\{\bar{\pi}_{u}:u\in\Delta(\mathcal{B})\right\}$ as $\mathcal{B}$’s each-step mixing policy class.
An each-step mixing policy $\bar{\pi}_{u}\in\bar{\Pi}_{\mathcal{B}}$ can be executed by drawing $\pi\sim u$ freshly-at-random at each step $h\in[H]$ and taking action $a_{h}\sim\pi(\cdot|s_{h})$ (e.g. Li and Zhang (2022, 2023)).
@@
In offline imitation learning, the agent is given a collection of expert trajectories $\mathcal{D}=\left\{\tau_{1},\ldots,\tau_{N_{\mathrm{off}}}\right\}$, where $\tau_{i}=(s_{i,h},a_{i,h})_{h=1}^{H}$ is the $i$-th (reward-free) trajectory, all of which are drawn iid from the trajectory distribution of expert policy $\pi^{\mathrm{E}}$. Behavior Cloning finds policy $\pi\in\mathcal{B}$ that minimizes its log loss on expert’s actions on the seen states:
@@
Recent work of Foster et al. (2024) establishes a sharp horizon-independent sample complexity bound of Behavior Cloning, which we recall here:
Theorem 2 (Foster et al. (2024)).
@@
In interactive IL, the learner has the ability to query the demonstration expert interactively. A first way to model interaction with expert is through a trajectory-wise demonstration oracle $\mathcal{O}^{\mathrm{Traj}}$ Ross et al. (2011); Foster et al. (2024): given a state sequence $(s_{h})_{h=1}^{H}$, return $(a_{h})_{h=1}^{H}$ such that $a_{h}=\pi^{\mathrm{E}}(s_{h})$ for all $h$. Subsequent works have considered modeling the interaction with expert as interacting with a state-wise demonstration oracle Judah et al. (2014); Brantley et al. (2019); Nguyen and Daumé III (2020); Sekhari et al. (2024) $\mathcal{O}^{\mathrm{State}}$: given a state $s_{h}$ and step $h$, return $a_{h}=\pi^{\mathrm{E}}(s_{h})$. We consider the learner interacting with the environment and demonstration oracles using the following protocol:
@@
In practice, we expect the cost of querying $\mathcal{O}^{\mathrm{Traj}}$ to be higher than that of collecting a single offline expert trajectory Laskey et al. (2017a). Since $H$ queries to $\mathcal{O}^{\mathrm{State}}$ can simulate one query to $\mathcal{O}^{\mathrm{Traj}}$, the cost of a single $\mathcal{O}^{\mathrm{State}}$ query should be at least $\frac{1}{H}$ the cost of $\mathcal{O}^{\mathrm{Traj}}$. Consequently, we also expect one $\mathcal{O}^{\mathrm{State}}$ query to be more expensive than obtaining an additional offline (state, expert action) pair. We denote the ratio between these two costs as $C$, where $C\geq 1$ is an application-dependent constant.  [^2] For practical settings such as human-in-the-loop learning with expert interventions Mandlekar et al. (2020); Spencer et al. (2020), obtaining a short segment of corrective demonstrations may be cheaper than querying $\mathcal{O}^{\mathrm{State}}$ for each state therein. Thus, our cost model may need refinements to match practical applications, which we leave as interesting future work.
@@ 3 State-wise Annotation in Interactive Imitation Learning
Recent work Foster et al. (2024) on refined analysis of Behavior Cloning (BC) casts doubt on the utility of interaction in imitation learning: when measuring sample complexity in the number of trajectories annotated, BC is minimax optimal even among interactive algorithms (Foster et al., 2024, Corollary 2.1 and Theorem 2.2). Although benefits of interactive approaches have been shown in specific examples, progresses so far have been sparse Foster et al. (2024); Rajaraman et al. (2021), with the most general results in the less-practical tabular setting Rajaraman et al. (2021). In this section, we reexamine this conclusion and show that interaction benefits imitation learning in a fairly general sense in the general function approximation setting: when measuring sample complexity using the number of state-wise annotations, we design an interactive algorithm with sample complexity better than BC, as long as the expert policy has a low recovering cost $\mu$ in the environment.
@@
1:Input: MDP $\mathcal{M}$, state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$ with query budget $N_{\mathrm{int}}$, stationary policy class $\mathcal{B}$, online learning oracle $\mathbb{A}$.
@@ 3.1 Interactive IL Enables Improved Sample Complexity with State-wise Annotations
Our algorithm Stagger (short for State-wise DAgger), namely Algorithm 1, interacts with the demonstration expert using a state-wise annotation oracle $\mathcal{O}^{\mathrm{State}}$. Similar to the original DAgger Ross et al. (2011), it requires base policy class $\mathcal{B}$ and reduces interactive imitation learning to no-regret online learning. At iteration $n$, it rolls out the current policy $\pi^{n}$ obtained from an online learning oracle $\mathbb{A}$ and samples state $s^{n}$ from $d^{\pi^{n}}$. A classical example of $\mathbb{A}$ is the exponential weight algorithm that chooses policies from $\bar{\Pi}_{\mathcal{B}}$ (Cesa-Bianchi and Lugosi (2006); see also Proposition 41 in Appendix F). It then queries $\mathcal{O}^{\mathrm{State}}$ to get expert action $a^{*,n}$ and updates $\mathbb{A}$ with loss function $\ell^{n}(\pi)$ induced by this new example (Eq. (1)). The final policy $\hat{\pi}$ is returned as a uniform first-step mixture of the historical policies $\{\pi^{n}\}_{n=1}^{N_{\mathrm{int}}}$, i.e., sample one $\pi^{n}$ uniformly at random and execute it for the episode. In contrast to the DAgger variants analyzed in Foster et al. (2024); Rajaraman et al. (2021) , which trains a distinct policy at each step—yielding $H$ policies in total—and employs trajectory-level annotations, our algorithm utilizes parameter sharing of the policy representation across all steps and uses state-wise annotations. [^3] A one-sample-per-iteration version of AggreVate Ross and Bagnell (2014) has been analyzed in (Agarwal et al., 2019, Section 15.5). Our analysis follows a similar structure and further takes direct advantage of the expert action feedback in DAgger and the deterministic realizability assumption to get refined sample complexity.
@@
Suppose Stagger is run with a state-wise expert annotation oracle $\mathcal{O}^{\mathrm{State}}$, an MDP $\mathcal{M}$ where $(\mathcal{M},\pi^{\mathrm{E}})$ is $\mu$-recoverable, a policy class $\mathcal{B}$ such that deterministic realizability (Assumption 1) holds, and the online learning oracle $\mathbb{A}$ set as the exponential weight algorithm with decision space $\bar{\Pi}_{\mathcal{B}}$. Then it returns $\hat{\pi}$ such that, with probability at least $1-\delta$,
@@
Theorem 3 shows that Stagger returns a policy of suboptimality $O(\frac{\mu H\log B}{N_{\mathrm{int}}})$ using $N_{\mathrm{int}}$ interactive state-wise annotations from the expert. In comparison, with the cost of $N_{\mathrm{int}}$ state-wise annotations, one can obtain $\frac{CN_{\mathrm{int}}}{H}$ trajectory-wise annotations; Foster et al. (2024)’s analysis shows that Behavior Cloning with this number of trajectories from $\pi^{\mathrm{E}}$ returns a policy of suboptimality $O(\frac{RH\log B}{CN_{\mathrm{int}}})$ (recall Theorem 2). Thus, if $C\ll\frac{R}{\mu}$, Algorithm 1 has a better cost-efficiency guarantee than Behavior Cloning.
We now sketch the proof of Theorem 3. In line with Foster et al. (2024), we define the online, on-policy state-wise estimation error as
@@
The proof proceeds by bounding this error and translating it to the performance difference between $\hat{\pi}$ and $\pi^{\mathrm{E}}$. While our definition of estimation error is similar to (Foster et al., 2024, Appendix C.2), their definition requires all $H$ states per trajectory, while ours depends on the distribution over a state sampled uniformly from the rollout of policy $\pi^{n}$. This enables each labeled state to serve as immediate online feedback, fully utilizing the adaptivity of online learning.